OpenStudy (samara8954):

Which equations show that the set of whole numbers is not closed under subtraction? Choose all answers that are correct. A. 1 – (–2) = 3 B. 1 – 2 = –1 C. 2 – 0 = 2 D. 2 – 4 = –2

OpenStudy (samara8954):

@water_girl_101 @freckles @Mimi_x3

OpenStudy (samara8954):

can someone help me?

OpenStudy (mimi_x3):

Okk whole numbers are integers that are 0 and greater so its 0,1,2,3,4,5 ......

OpenStudy (mimi_x3):

So which equations when you subtract do you NOT get a whole number or rather do you get a negative number?

OpenStudy (samara8954):

a and c

OpenStudy (mimi_x3):

Ok so A and C we subtracted and still have a whole number or rather we still have a positive number We wanna find an equation where we subtract 2 whole numbers and end up with a negative integer Additionally If you take a look in Option A we have 1 – (–2) = 3 And -2 is a negative integer and hence not even a whole number

OpenStudy (samara8954):

so a and c is the correct answer

OpenStudy (mimi_x3):

ughhh Nope cuz its still a positive integer

OpenStudy (mimi_x3):

Ok lemme try to explain again

OpenStudy (samara8954):

so b and d

OpenStudy (mimi_x3):

yes lol

OpenStudy (samara8954):

Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (samara8954):

can someone help me with this one

OpenStudy (mimi_x3):

okkk ummm I gotta run soon but i want you to attempt to solve it on your own

OpenStudy (mimi_x3):

So closed under addition means If you added the values in each set then you still will get a value within that set So lets take the set {0} If you add 0+0=0 Since 0 is in the set then we can conclude that the set {0} is closed under addition

OpenStudy (samara8954):

i think it is a d and b

OpenStudy (mimi_x3):

Lets look at the set " all even integers" So an even number is represented by 2k where k is an integer so for instance when k=1 we have 2*1=2 When k=2 we have 2*2=4 so basically 2k represents all even functions so 2k+2k=4k=2(2k) so we can conclude that all even integers are closed under addition

OpenStudy (samara8954):

am I corrrect about a d and b

OpenStudy (samara8954):

yes I am

OpenStudy (mimi_x3):

ok good

OpenStudy (mimi_x3):

Ok so I dont have time to explain rational numbers cuz I am running out but a rational number is closed under addition

OpenStudy (samara8954):

so its not a d and b

OpenStudy (mimi_x3):

Lemme get @freckles to help you

OpenStudy (samara8954):

@freckles it offfline

OpenStudy (mimi_x3):

ok so basically a is the only one that isnt closed under addition The rest are

OpenStudy (freckles):

hello

OpenStudy (freckles):

is this the question you are looking at: "Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}"

OpenStudy (freckles):

A) we can look at all the possibilities or we can try to find a counterexample to it being closed...(you know something showing that it is not false) or let's look at all the possibilities when you add any of the following, do you get a number inside the set {0,2,5,8} always: 0+0=? 0+2=? 0+5=? 0+8=? 2+2=? 2+0=? 2+5=? 2+8=? 5+0=? 5+2=? 5+5=? 5+8=? 8+0=? 8+2=? 8+5=? 8+8=? as soon as you run into a sum that is not an element of {0,2,5,8} you know the set {0,2,5,8} will not be closed under addition so what do you think about {0,2,5,8} closed or not closed under addition?

OpenStudy (samara8954):

i think the answer is even integers and rational numbers

OpenStudy (freckles):

what about the set just containing 0 0+0 's sum is an element of {0} is it not?

OpenStudy (samara8954):

or it could be just a

OpenStudy (freckles):

why do you think a?

OpenStudy (samara8954):

so it is b and c

OpenStudy (freckles):

think about the sums I mentioned above for example is 2+8 an element {0,2,5,8}

OpenStudy (samara8954):

no

OpenStudy (freckles):

so A is definitely not closed under addition

OpenStudy (freckles):

let's think about the even integers 2a is even for any integer a 2i is even for any integer i what about the sum of 2a and 2i 2a+2i=2(a+i) what does this say a bout the set of even integers under addition ?

OpenStudy (samara8954):

yes

OpenStudy (freckles):

yes the set of even integers is totally closed under addition

OpenStudy (freckles):

now let's move on to the rationals m/n and c/b are elements of the rationals (Assuming of course n and b are not zero you know since we can not divide by 0 and also m,n,c,b are integers ) so what happens when we add m/n and c/b

OpenStudy (freckles):

\[\frac{m}{n}+\frac{c}{b}=\frac{m}{n} \cdot \frac{b}{b}+\frac{c}{b} \cdot \frac{n}{n}=\frac{mb+cn}{nb}\]

OpenStudy (freckles):

if m,n,c,b are integers then is mb+cn an integer?

OpenStudy (freckles):

I will give you a hint the set of integers is closed under addition and multiplication

OpenStudy (samara8954):

so it is c and d

OpenStudy (samara8954):

or b and c

OpenStudy (freckles):

I thought you agreed with me that the set of even integers was closed under addition we are now talking about the rationals an element of the rationals looks like m/n or you can say c/b where m,n,c,b are integers (and n and b are not 0) we are adding m/n and c/b to see if the set of rationals is closed under addition the rationals are numbers that can be written as a fraction where the numerator and denominator (not 0) are integer numbers

OpenStudy (freckles):

that is why I asked you to tell me if the sum of m/n and c/b 's numerator was an integer then I was going to ask about the denominator

OpenStudy (freckles):

\[\frac{m}{n}+\frac{c}{b}=\frac{mb+cn}{nb}\] I gave you the hint that the set of integers is closed under addition and multiplication which means both mb+cn and nb are integers which means m/n+c/b is an element of the rationals which means the set of rationals is closed under addition

OpenStudy (freckles):

and we already discussed the last choice {0} 0+0 is definitely an element of {0}

OpenStudy (samara8954):

so b and c are correct or is it b and d

OpenStudy (freckles):

I'm confused. Which of my explanations are you having trouble with?

OpenStudy (freckles):

And why do you assume you have only 2 choices of the 4 that are correct

OpenStudy (freckles):

again 2a+2i=2(a+i) which is still an even number and m/n+c/b=(mb+nc)/nb where both mb+nc and nb are integers (where nb is not zero) so m/n+c/b is still rational 0+0 is an element of 0 and finally 2+8 is not an element of {0,2,5,8}

OpenStudy (freckles):

is an element of {0}

OpenStudy (samara8954):

so it it is b c and d

OpenStudy (freckles):

yes do you understand why ?

OpenStudy (freckles):

if not can you please tell me which of my explanations doesn't make sense?

OpenStudy (samara8954):

yes

OpenStudy (samara8954):

Numbers p and q are negative integers. Which statements are always true? Choose all answers that are correct. A. p + q is a negative integer B. p • q is a positive integer C. p – q is a negative integer D. is a negative integer

OpenStudy (samara8954):

i think this one is a and c

OpenStudy (freckles):

so a negative number times a negative number isn't positive ?

OpenStudy (freckles):

also what is -5-(-5)?

OpenStudy (freckles):

and I can't tell what D is

OpenStudy (samara8954):

so is it a and c

OpenStudy (anonymous):

yep its a and c